# Bessel potentials and Green functions on pseudo-Euclidean spaces

### v1

@@ 5.2 Green functions of the Klein Gordon equation
The massive and massless cases are quite similar and they often appear in physical applications. The tachyonic case is more exotic, but also interesting.
@@
$G^{\mathrm{F}/\overline{\mathrm{F}}}(x)=\int\frac{{\rm e}^{{\rm i}px}}{(-E+p^{2}\mp{\rm i}0)}\frac{{\rm d}p}{(2\pi)^{d}}.$ (5.22)
As follows from a general theory of hyperbolic equations, the Klein-Gordon equation possesses also another important pair of Green functions: the retarded (or forward) Green function $G^{\vee}$ and the advanced (backward) Green function $G^{\wedge}$. They are uniquely defined by the conditions
@@
We will also consider distinguished solutions of the (homogeneous) Klein-Gordon equation, that is functions $G^{\circ}$ satisfying
@@ 5.3 Massive Klein-Gordon equation
Specifying Theorem 5.2 to $\mu=2$, we obtain the following expressions for the Feynman and anti-Feynman Green functions:
Theorem 5.3
@@
The retarded and advanced Green functions of the Klein-Gordon equation are obtained by specifying Theorem 5.2 to $\mu=2$. In the following theorem, we also identify their regular and singular part.
Theorem 5.4
@@
Proof.  The formula for the Green function of the Klein-Gordon equation, denoted as $G^{\vee/\wedge}$, is given by equation (5.32), which was computed earlier for general $\mu$ (5.7). The decomposition into (5.34) and (5.35) is due to (2.48). For even $d$, the decomposition can be rewritten using (2.49). $\Box$
Introduce (5.38), (5.40) distinguished solutions of the Klein–Gordon equation $-\Box+m^{2}$.
@@
We will call distinguished Green functions and solutions jointly “two-point functions”. $G_{m}^{\mathrm{PJ}}$ is supported in $J^{\vee}\cup J^{\wedge}$. Here are the expressions for these solutions in terms of positions:
@@
Note the identities satisfied by the two-point functions:
@@ 5.4 Massless Klein-Gordon equation
The massless case is quite similar to the massive one: we need only to set $m=0$ in the previous subsection. In particular, all identities (5.44) are satisfied. There are a few simplifications. Only the most singular part of the massive propagator remains in the massless case.
Theorem 5.5
@@ 5.5 Tachyonic Klein-Gordon equation
We have a minor notational problem how to indicate that we replaced $m^{2}$ with $-m^{2}$. Putting $+{\rm i}m$ or $-{\rm i}m$ is not appropriate, because it means the analytic interpretation of the corresponding function in the mass parameter. Therefore, in the following theorem we leave blank the place for a subscript.
@@
Theorem 5.6
$G^{\mathrm{F}}(x)/G^{\overline{\mathrm{F}}}(x)$ $=\int\frac{{\rm e}^{{\rm i}px}}{(-m^{2}+p^{2}\mp{\rm i}0)}\frac{{\rm d}p}{(2\pi)^{d}}$ (5.54)
@@
$G^{\mathrm{F}}(x)/G^{\overline{\mathrm{F}}}(x)=\frac{\pm{\rm i}\sqrt{\pi}m^{d-2}{\rm e}^{\mp{\rm i}\pi(\frac{d}{2}-1)}}{(4\pi)^{\frac{d}{2}}}U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}(-x^{2}\mp{\rm i}0)}{4}\Big{)}$
@@
$G^{\mathrm{F}}(x)/G^{\overline{\mathrm{F}}}(x)=\frac{\pm{\rm i}\sqrt{\pi}m^{d-2}{\rm e}^{\mp{\rm i}\pi(\frac{d}{2}-1)}}{(4\pi)^{\frac{d}{2}}}U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}|x^{2}|}{4}\Big{)}.$
@@
$G^{\vee/\wedge}$ are not tempered distributions on ${\mathbb{R}}^{1,d-1}$, and therefore they cannot be expressed in terms of the Fourier transformation in all variables, as in the massive and massless cases. However, they are well-defined, and in the following theorem we give three equivalent formulas for these propagators.
Theorem 5.7
@@
They are supported in $J^{\vee}$, resp. $J^{\wedge}$. We can decompose $G^{\vee/\wedge}$ into a singular and regular part:
$G^{\vee/\wedge}(x)$ $=G_{\rm sing}^{\vee/\wedge}(x)+G^{\vee/\wedge}_{\text{reg}}(x).$ (5.61)
@@
$G_{\rm sing}^{\vee/\wedge}(x)$ $=\frac{\theta(\pm x^{0})}{2\pi^{\frac{d}{2}-1}}\sum_{j=0}^{\frac{d-5}{2}}\frac{1}{j!\Gamma(2-\frac{d}{2}+j)}\Big{(}\frac{m^{2}}{4}\Big{)}^{j}(x^{2})_{-}^{1-\frac{d}{2}+j},$ (5.62)
$G_{\rm reg}^{\vee/\wedge}(x)$ $=\frac{\theta(\pm x^{0})}{2\pi^{\frac{d}{2}-1}}\sum_{j=\frac{d-3}{2}}^{\infty}\frac{1}{j!\Gamma(2-\frac{d}{2}+j)}\Big{(}\frac{m^{2}}{4}\Big{)}^{j}(-x^{2})^{1-\frac{d}{2}+j}\theta(-x^{2}).$ (5.63)
@@
$G_{\rm sing}^{\vee/\wedge}(x)$ $=\theta(\pm x^{0})\frac{1}{2\pi^{\frac{d}{2}-1}}\sum_{j=0}^{\frac{d}{2}-2}\frac{1}{(\frac{d}{2}-2-j)!}\biggl{(}\frac{m^{2}}{4}\biggr{)}^{\frac{d}{2}-2-j}\delta^{(j)}(x^{2}),$ (5.64)
$G_{\rm reg}^{\vee/\wedge}(x)$ $=\theta(\pm x^{0})\frac{2\pi m^{d-2}}{(4\pi)^{\frac{d}{2}}}{\bf F}_{\frac{d}{2}-1}\Big{(}\frac{m^{2}|x^{2}|}{4}\Big{)}\theta(-x^{2}).$ (5.65)
@@
Let us compute the sum of the tachyonic $G^{\mathrm{F}}(x)$ and $G^{\overline{\mathrm{F}}}(x)$:
$G^{\mathrm{F}}(x)+G^{\overline{\mathrm{F}}}(x)$ $=\frac{2\pi m^{d-2}}{(4\pi)^{\frac{d}{2}}}{\bf F}_{\frac{d}{2}-1}\Big{(}-\frac{m^{2}x^{2}}{4}\Big{)},\quad x^{2}>0;$ (5.67)
$G^{\mathrm{F}}(x)+G^{\overline{\mathrm{F}}}(x)$ $=\frac{4\pi}{(4\pi)^{\frac{d}{2}}}\Big{(}\frac{-x^{2}}{4}\Big{)}^{1-\frac{d}{2}}{\bf F}_{1-\frac{d}{2}}\Big{(}-\frac{m^{2}x^{2}}{4}\Big{)},\quad x^{2}<0.$ (5.68)
Thus $G^{\mathrm{F}}(x)+G^{\overline{\mathrm{F}}}(x)$ does not have a causal support, and consequently,
$G^{\mathrm{F}}(x)+G^{\overline{\mathrm{F}}}(x)\neq G^{\vee}(x)+G^{\wedge}(x).$ (5.69)
@@
In the tachyonic case, we do not have the solutions $G^{(\pm)}$. However, we can define the Pauli-Jordan propagator
$G^{\mathrm{PJ}}(x)$ $=\frac{1}{(2\pi)^{d-1}}\int{\rm d}\vec{p}\,{\rm e}^{{\rm i}\vec{x}\vec{p}}\frac{\sin\left(x^{0}\sqrt{\vec{p}^{2}-m^{2}}\right)}{\sqrt{\vec{p}^{2}-m^{2}}}$ (5.70)
@@
Note that $G^{\mathrm{PJ}}$ cannot be written in the form (5.25).
@@ 5.6 Averages of plane waves on the hyperbolic plane
The Minkowski space possesses two kinds of hyperboloids. The two-sheeted hyperboloid consists of two connected components isomorphic to the hyperbolic space. In this subsection, we compute the Fourier transform of the natural measure on one of the sheets of the two-sheeted hyperboloids, similarly as in the Theorem 3.3.
Consider the future/past hyperboloid in $d$-dimensional Minkowski space, denoted ±${\mathbb{H}}_{\pm,m}={\mathbb{H}}_{\pm,m}^{d-1}$, consisting of points $p$ such that $p^{2}+m^{2}=0$ and ±$\pm p^{0}>0$. Let ${\rm d}\Omega_{m}$ denote the standard measure on ±,msubscriptplus-or-${\mathbb{H}}_{\pm,m}$. We will see that up to a coefficient its Fourier transform is essentially the “positive frequency solution of the Klein-Gordon equation.”
Theorem 5.8
@@
Proof.  This average, up to a coefficient, coincides with $G^{(\pm)}$ defined in (5.40), which we have already computed:
@@ 5.7 Averages of plane waves on the deSitter space
$=\frac{m(2\pi)^{d}}{\pi{\rm i}}\big{(}G^{\mathrm{F}}(x)-G^{\overline{\mathrm{F}}}(x)\big{)}.$ (5.79)
@@ Acknowledgement.
The support of the National Science Center of Poland under the grant UMO-2019/35/B/ST1/01651 is acknowledged.

### v2

@@ 5.2 Green functions of the Klein Gordon equation
The massive and massless cases are quite similar and they often appear in physical applications. They are often discussed in detail in the literature. The tachyonic case is more exotic and less known, but also interesting.
@@
$G_{m}^{\mathrm{F}/\overline{\mathrm{F}}}(x)=\int\frac{{\rm e}^{{\rm i}px}}{(-E+p^{2}\mp{\rm i}0)}\frac{{\rm d}p}{(2\pi)^{d}}.$ (5.22)
As follows from a general theory of hyperbolic equations, the Klein-Gordon equation (5.17) possesses also another important pair of Green functions: the retarded (or forward) Green function $G^{\vee}$ and the advanced (backward) Green function $G^{\wedge}$. They are uniquely defined by the conditions
@@
Note that the above definition provides $G^{\vee/\wedge}$ for all $E\in{\mathbb{R}}$. In the case $-E\geq 0$, with $-E=m^{2}$ they coincide with $G_{m}^{\vee/\wedge}$ defined already with the help of Fourier transformation. In the tachyonic case they will be denoted $G_{{\rm i}m}^{\vee/\wedge}=G_{-{\rm i}m}^{\vee/\wedge}$ and they need a separate discussion, see Subsection 5.5
We will also consider certain distinguished solutions of the (homogeneous) Klein-Gordon equation, that is functions $G^{\circ}$ satisfying
@@ 5.3 Massive Klein-Gordon equation
Specifying Theorem 10 to $\mu=2$, we obtain the following expressions for the Feynman and anti-Feynman Green functions:
Theorem 12
@@
The retarded and advanced Green functions of the Klein-Gordon equation are obtained by specifying Theorem 10 to $\mu=2$. In the following theorem, we also identify their regular and singular part.
Theorem 13
@@
Proof.  The formula for the Green function of the Klein-Gordon equationis given by equation (5.32), which was computed earlier for a general $\mu$ (5.7). The decomposition into (5.34) and (5.35) is due to (2.48). For even $d$, the decomposition can be rewritten using (2.49). $\Box$
Introduce the following distinguished solutions of the Klein–Gordon equation $-\Box+m^{2}$:
@@
Following [13], we will call distinguished Green functions and solutions jointly propagators. $G_{m}^{\mathrm{PJ}}$ is supported in $J^{\vee}\cup J^{\wedge}$. Here are the expressions for these solutions in terms of positions:
@@
Note the identities satisfied by the propagators:
@@ 5.4 Massless Klein-Gordon equation
The massless case is quite similar to the massive one: we need only to set $m=0$ in the previous subsection. In particular, all identities (5.44) are satisfied. There are a few simplifications. Only the most singular part of the massive propagator remains in the massless case. This is the special case of Riesz potentials, massless limit of Bessel potentials, that we studied in the section 3.2
Theorem 14
@@ 5.5 Tachyonic Klein-Gordon equation
Usually, tachyonic quantum fields are considered to be unphysical [21]. Nevertheless, every now and then there are attempts to analyze them in the physics literature, see [22], and more recently [23].
We have a minor notational problem how to indicate that we replaced $m^{2}$ with $-m^{2}$. Naively, one would think it should be indicated by both $+{\rm i}m$ and $-{\rm i}m$ instead of $m$. However, this would suggest the analytic continuation ${\rm e}^{{\rm i}\phi},$$\pm\phi\in[0,\pi]$, which is not always appropriate. This problem appears in the case of the Feynman propagator: we will write $G_{{\rm i}m}^{\mathrm{F}}$, but not $G_{-{\rm i}m}^{\mathrm{F}}$. Similarly, for the anti-Feynman propagator we will write $G_{-{\rm i}m}^{\overline{\mathrm{F}}}$, but not $G_{{\rm i}m}^{\overline{\mathrm{F}}}$. In the case of retarded/advanced propagators, this problem will be absent, since the analytic continuation can be performed in $m^{2}$: thus $G_{{\rm i}m}^{\wedge/\vee}=G_{-{\rm i}m}^{\wedge/\vee}$.
@@
Theorem 15
$G_{{\rm i}m}^{\mathrm{F}}(x)/G_{-{\rm i}m}^{\overline{\mathrm{F}}}(x)$ $=\int\frac{{\rm e}^{{\rm i}px}}{(-m^{2}+p^{2}\mp{\rm i}0)}\frac{{\rm d}p}{(2\pi)^{d}}$ (5.54)
@@
$G_{{\rm i}m}^{\mathrm{F}}(x)/G_{-{\rm i}m}^{\overline{\mathrm{F}}}(x)=\frac{\pm{\rm i}\sqrt{\pi}m^{d-2}{\rm e}^{\mp{\rm i}\pi(\frac{d}{2}-1)}}{(4\pi)^{\frac{d}{2}}}U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}(-x^{2}\mp{\rm i}0)}{4}\Big{)}$
@@
$G_{{\rm i}m}^{\mathrm{F}}(x)/G_{-{\rm i}m}^{\overline{\mathrm{F}}}(x)=\frac{\pm{\rm i}\sqrt{\pi}m^{d-2}{\rm e}^{\mp{\rm i}\pi(\frac{d}{2}-1)}}{(4\pi)^{\frac{d}{2}}}U_{\frac{d}{2}-1}\Big{(}\frac{m^{2}|x^{2}|}{4}\Big{)}.$
@@
Unfortunately, the tachyonic Feynman and antiFeynman propagator do not have the usual physical interpretation, as the vacuum expectation value of the time-ordered, resp. anti-time-ordered product of fields. In fact, for tachyons the vacuum is ill defined. Nevertheless, some authors, e.g. [22], try to use the above Feynman propagator to define interacting tachyonic quantum field theory.
Retarded and advanced tachyonic Green functions $G_{{\rm i}m}^{\vee/\wedge}$ are not tempered distributions on ${\mathbb{R}}^{1,d-1}$, and therefore they cannot be expressed in terms of the Fourier transformation in all variables, as in the massive and massless cases (5.30). However, they are well-defined, and in the following theorem we give three equivalent formulas for these propagators.
Theorem 16
@@
They are supported in $J^{\vee}$, resp. $J^{\wedge}$. We can decompose $G_{{\rm i}m}^{\vee/\wedge}$ into a singular and regular part:
$G_{{\rm i}m}^{\vee/\wedge}(x)$ $=G_{{\rm i}m,{\rm sing}}^{\vee/\wedge}(x)+G^{\vee/\wedge}_{{\rm i}m,\text{reg}}(x).$ (5.61)
@@
$G_{{\rm i}m,{\rm sing}}^{\vee/\wedge}(x)$ $=\frac{\theta(\pm x^{0})}{2\pi^{\frac{d}{2}-1}}\sum_{j=0}^{\frac{d-5}{2}}\frac{1}{j!\Gamma(2-\frac{d}{2}+j)}\Big{(}\frac{m^{2}}{4}\Big{)}^{j}(x^{2})_{-}^{1-\frac{d}{2}+j},$ (5.62)
$G_{{\rm i}m,{\rm reg}}^{\vee/\wedge}(x)$ $=\frac{\theta(\pm x^{0})}{2\pi^{\frac{d}{2}-1}}\sum_{j=\frac{d-3}{2}}^{\infty}\frac{1}{j!\Gamma(2-\frac{d}{2}+j)}\Big{(}\frac{m^{2}}{4}\Big{)}^{j}(-x^{2})^{1-\frac{d}{2}+j}\theta(-x^{2}).$ (5.63)
@@
$G_{{\rm i}m,{\rm sing}}^{\vee/\wedge}(x)$ $=\theta(\pm x^{0})\frac{1}{2\pi^{\frac{d}{2}-1}}\sum_{j=0}^{\frac{d}{2}-2}\frac{1}{(\frac{d}{2}-2-j)!}\biggl{(}\frac{m^{2}}{4}\biggr{)}^{\frac{d}{2}-2-j}\delta^{(j)}(x^{2}),$ (5.64)
$G_{{\rm i}m,{\rm reg}}^{\vee/\wedge}(x)$ $=\theta(\pm x^{0})\frac{2\pi m^{d-2}}{(4\pi)^{\frac{d}{2}}}{\bf F}_{\frac{d}{2}-1}\Big{(}\frac{m^{2}|x^{2}|}{4}\Big{)}\theta(-x^{2}).$ (5.65)
@@
Let us compute the sum of the tachyonic Feynman and antiFeynman propagator:
$G_{{\rm i}m}^{\mathrm{F}}(x)+G_{-{\rm i}m}^{\overline{\mathrm{F}}}(x)$ =$=\begin{cases}\frac{2\pi m^{d-2}}{(4\pi)^{\frac{d}{2}}}{\bf F}_{\frac{d}{2}-1}\Big{(}-\frac{m^{2}x^{2}}{4}\Big{)},&\quad x^{2}>0;\\ \frac{4\pi}{(4\pi)^{\frac{d}{2}}}\Big{(}\frac{-x^{2}}{4}\Big{)}^{1-\frac{d}{2}}{\bf F}_{1-\frac{d}{2}}\Big{(}-\frac{m^{2}x^{2}}{4}\Big{)},&\quad x^{2}<0.\end{cases}$ (5.67)
Thus $G_{{\rm i}m}^{\mathrm{F}}(x)+G_{-{\rm i}m}^{\overline{\mathrm{F}}}(x)$ does not have a causal support, and consequently,
$G_{{\rm i}m}^{\mathrm{F}}(x)+G_{-{\rm i}m}^{\overline{\mathrm{F}}}(x)\neq G_{{\rm i}m}^{\vee}(x)+G_{{\rm i}m}^{\wedge}(x).$ (5.68)
@@
In the tachyonic case, we do not have the solutions $G_{{\rm i}m}^{(\pm)}$. However, we can define the Pauli-Jordan propagator
$G_{{\rm i}m}^{\mathrm{PJ}}(x)=G_{-{\rm i}m}^{\mathrm{PJ}}$ $=\frac{1}{(2\pi)^{d-1}}\int{\rm d}\vec{p}\,{\rm e}^{{\rm i}\vec{x}\vec{p}}\frac{\sin\left(x^{0}\sqrt{\vec{p}^{2}-m^{2}}\right)}{\sqrt{\vec{p}^{2}-m^{2}}}$ (5.69)
@@
Note that $G_{{\rm i}m}^{\mathrm{PJ}}$ cannot be written in the form (5.25).
@@ 5.6 Averages of plane waves on the hyperbolic plane
The Minkowski space possesses two kinds of hyperboloids. The two-sheeted hyperboloid consists of two connected components isomorphic to the hyperbolic space. In this subsection, we compute the Fourier transform of the natural measure on one of the sheets of the two-sheeted hyperboloids, similarly as in the Theorem 4.
Consider the future/past hyperboloid in the $d$-dimensional Minkowski space, denoted ±${\mathbb{H}}_{\pm,m}={\mathbb{H}}_{\pm,m}^{d-1}$, consisting of points $p$ such that $p^{2}+m^{2}=0$ and ±$\pm p^{0}>0$. Let ${\rm d}\Omega_{m}$ denote the standard measure on ±,msubscriptplus-or-${\mathbb{H}}_{\pm,m}$. We will see that up to a coefficient its Fourier transform is essentially the “positive frequency solution of the Klein-Gordon equation.”
Theorem 17
@@
Proof.  This average, up to a coefficient, coincides with $G_{m}^{(\pm)}$ defined in (5.40), which we have already computed:
@@ 5.7 Averages of plane waves on the deSitter space
$=\frac{m(2\pi)^{d}}{\pi{\rm i}}\big{(}G_{{\rm i}m}^{\mathrm{F}}(x)-G_{-{\rm i}m}^{\overline{\mathrm{F}}}(x)\big{)}.$ (5.78)
@@ Acknowledgement.
The support of the National Science Center of Poland under the grant UMO-2019/35/B/ST1/01651 is acknowledged. We thank the referees for their remarks. In particular, we are grateful to one of them for drawing our attention to Hörmander’s criterion for multiplication of distribution; see its application at the end of Subsection 5.1